# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v1

@@ V Experimental Evaluation
In this section, we first analyze the performance of the proposed HFLDD and compare it with baseline methods to demonstrate how our method can mitigate the impact of quantity-based label distribution skewness. Subsequently, we compare the communication overhead of different methods. Finally, we evaluate the impact of $K$ in the K-Means algorithm on the performance of HFLDD.
V-A Experimental Setup
V-A1 Datasets
@@
Our experiments are conducted on two widely used datasets MNIST [46], and CIFAR10 [47]. MNIST consists of 50,000 binary images of handwritten digits, which can be divided into 10 classes. CIFAR10 consists of 50,000 daily object images, which can also be divided into 10 classes. In our Non-IID setting, we investigate scenarios where each client possesses 1, 2, or 10 of the ten classes. In our experiments, each dataset is divided into 80% for training and 20% for testing.
@@
Apart from the training and testing local client dataset, we also need additional global datasets to generate soft labels for label knowledge collection. We randomly select 1,000 samples from other similar datasets and construct the corresponding global dataset. For MNIST experiment, the global dataset is constructed from the FMNIST dataset [48]. For CIFAR10 experiments, the global dataset is constructed from the Tinyimagenet dataset [49].
@@
Figure 3: The performance of the global model on MNIST and CIFAR10. It can be observed that in cases with severe Non-IID setting, the global model of HFLDD still converges rapidly to satisfactory test accuracy.
V-A2 Models
In our experiments, the applied models are all based on Convolutional Neural Network (CNN) [1]. The model architectures for different datasets are depicted in Fig. 4. For the MNIST dataset, we construct a shallow CNN model suitable for handwritten digit classification. The total number of model parameters are 44,426. For the CIFAR10 dataset, we increase the depth of the model to better capture the features of complex images in CIFAR10. The total number of model parameters are 1,020,160. The input for both models is $32\times 32\times 3$ image and the output is a 10-dimensional one-hot class label. It should be noted that the proposed HFLDD is not limited by the model architectures, and it can be applied to any model. In the experiments, we only show the results based on the above mentioned models.
(a) MNIST
(b) CIFAR10
Figure 4: Model architecture for MNIST and CIFAR10 datasets.
@@ V-A2 Models
Figure 5: Heterogeneous topology formed by client clustering, where $N=100$, $K=10$. The bottom right corner of each subfigure shows the results of homogeneous clustering, where different colors represent clients from different homogeneous clusters. In each subfigure, the clients connected with the server are the clusters headers. Each cluster header and its connected clients construct a heterogeneous cluster.
V-A3 Baselines
To evaluate the performance of our proposed HFLDD algorithm, we select three typical baseline methods for comparison:
  • FedAvg [9], a classic FL algorithm, where each client conducts local training and the server aggregates all the local models for global model update;
  • FedSeq [32], a classic HFL algorithm, which utilizes client clustering and sequential training to enhance model performance;
@@ V-A4 Implementation Details
In our experiments, conducted on a machine equipped with a Tesla V100 GPU and implemented using PyTorch, we fixed the number of clients $N$ to 100. For the pre-training phase, both MNIST and CIFAR10 datasets were used with the following settings: learning rate $\eta=0.01$, local update rounds $E_{1}=10$, a batch size of 64, and the SGD optimizer. These settings ensure that each client can effectively extract knowledge from its local data, thereby enhancing the discriminability for subsequent clustering tasks. For constructing the heterogeneous topology, we set K-Means clustering with 10 clusters. Due to the need to transmit distilled images between clients, we set each client to distill $n=100$ images. The learning rate for distilled data $\eta_{k}=0.004$. Following the transmission of distilled data, formal optimization of the global model begins. The server performs updates for $T=300$ rounds, with $E_{2}=2$ local update rounds, a batch size of 32, and all other configurations consistent with the pre-training phase.
@@
To ensure a fair comparison, FedAvg and FedSeq are both configured with the same number of global update rounds and local update epochs as HFLDD. Additionally, in the FedSeq algorithm, the number of clusters are set as 10, with each cluster containing 10 clients. For centralized learning, the training batch size is 64, and the learning rate is 0.01.
V-B Training Performance
@@
To comprehensively show the training performance of the proposed HFLDD, we compare HFLDD with the baseline methods in both Non-IID and IID scenarios.
V-B1 Non-IID Scenario
@@
We consider two severe Non-IID settings: 1) $C=1$, where each client holds only one class of data, making it the most challenging setting; 2) $C=2$, where each client holds two classes of data. The data labels among different clients are still severely imbalanced but partially overlap.
@@
In Fig. 3, we compare the training performance of HFLDD with baseline methods on the datasets of MNIST and CIFAR10, respectively. When $C=1$, as shown in Fig. 3a and Fig. 3d, it can be observed that our proposed HFLDD algorithm outperforms FedAvg and FedSeq on both MNIST and CIFAR10. Specifically, on the CIFAR10 dataset, HFLDD achieves an accuracy of 64.3% after 300 communication rounds, far exceeding FedAvg at 27.0% and FedSeq at 19.6%. On the MNIST dataset, although the final accuracy of HFLDD is similar to FedSeq, it is clear that HFLDD converges faster. When $C=2$, on the CIFAR10 dataset, as shown in Fig. 3e, HFLDD achieves a final test accuracy of 70.1%, which also outperforms FedAvg and FedSeq. On the MNIST dataset, as shown in Fig. 3b, HFLDD and FedSeq achieve nearly identical training results, but both outperform FedAvg. Additionally, under the Non-IID settings described above, the performance gap between HFLDD and CL is the smallest, in comparison with FedAvg and FedSeq.
@@
Figs. 5a, 5b, 5d, and 5e specifically illustrate the heterogeneous topology formed by HFLDD in the Non-IID scenario in the experiments. Because of the significant differences in data distribution among the clients, the clustering algorithm can effectively classify the clients. In Fig. 5d, the number of clients in all heterogeneous clusters is the same. In Figs. 5a, 5b, and 5e, a few isolated nodes appear in the heterogeneous topology due to some clients being misclassified. However, they do not significantly affect overall performance. This is because the data distribution in most heterogeneous clusters is approximately the same as the global data distribution, effectively compensating for the overall performance.
Metrics HFLDD FedAvg FedSeq
Accuracy=80% rounds/Accuracy 5/80.63% 17/80.54% 1/84.32%
Traffic (MB) 39.92M 593.15M 38.98M
Ratio 1X 14.86X 0.9765X
Metrics HFLDD FedAvg FedSeq
Accuracy=70% rounds/Accuracy 32/70.00% 35/70.27% 4/71.16%
Traffic (MB) 5844.32M 27630.37M 2296.04M
Ratio 1X 4.7277X 0.3929X
V-B2 IID Scenario
In this case, we set $C=10$, where each client holds all ten classes of data and the data labels among all the clients are balanced. On the MNIST dataset, as shown in Fig. 3c, all the methods achieve almost identical test accuracy, and HFLDD’s convergence speed is intermediate among the baseline algorithms. On the CIFAR10 dataset, as shown in Fig. 3f, the baseline methods can achieve higher test accuracy than the proposed HFLDD. This can be attributed to two main factors. First, in HFLDD, the global model is trained based the dataset of the cluster headers and the distilled data from the cluster members. The information loss caused by dataset distillation degrades the training performance and makes the test accuracy of HFLDD lower than the baseline methods based on the original datasets under IID conditions. Secondly, as seen from Figs. 5c and 5f, when each client holds 10 classes of data, the clustering algorithms cannot distinguish the differences among clients, resulting in many imbalanced heterogeneous clusters. In this case, if the distilled images are transmitted among heterogeneous clusters, the cluster headers will have unbalanced sample quantities, resulting in a quantity skew[44], which breaks the original IID relationship among clients.
Nevertheless, compared to baseline distributed learning methods, the proposed HFLDD is more robust to Non-IID data. As the data distribution shifts from IID to Non-IID, i.e., from $C=10$ to $C=1$, as shown in Fig. 3, On the CIFAR10 dataset, the test accuracy of HFLDD drops from 72.2% to 64.3%. On the MINIST dataset, the test accuracy of HFLDD drops from 98.4% to 98.1%, demonstrating the robustness of HFLDD to Non-IID data. In contrast, the test accuracy of the baseline methods, FedSeq and FedAvg, rapidly declines or even fails to converge.
V-C Communication Cost
@@
Figure 6: Communication traffic consumption of three algorithms when achieving the target accuracy on MNIST and CIFAR10 under Non-IID conditions $(C=2)$
The analytical communication cost of HFLDD is provided in Secion. V-B. In this part, we will compare communication cost of HFLDD with the baseline distributed learning methods, FedAvg and FedSeq.
The communication cost of FedAvg in $T$ rounds of training is given by Equation 3.
For FedSeq, the communication cost comes from parameter sharing between clients within cluster and across clusters, it can be expressed as
@@
Considering that when the training converges, the communication rounds and test accuracy of different methods are different. it would be unfair to compare the communication cost under a given communication rounds. Therefore, we compare the number of communication rounds and total communication cost required by different algorithms to achieve a certain test accuracy.
In the scenario with IID data, we set the target test accuracy as 80.0% for the MNIST dataset. Table II shows the number of rounds and communication cost for each algorithm to reach the target test accuracy. The ratio denotes the multiple of the communication cost of each method relative to the HFLDD. It can be observed that the communication cost of HFLDD is slightly higher than that of FedSeq, and 14.86 times less than that of FedAvg. For the CIFAR10 dataset, we set the target accuracy as 70%. When each algorithm reaches the target test accuracy, the communication cost of HFLDD is still much smaller than that of FedAvg, while the communication cost consumed by FedSeq is 0.39 times that of HFLDD, as detailed in Table III. This is mainly because FedSeq converges faster than other algorithms when the data is IID.
@@
In the scenario with Non-IID data, when $C=1$, due to the severe imbalance of data labels, the baseline methods may not converge and it is infeasible to find proper target test accuracy that sufficiently reflects the communication cost of each algorithm. Therefore, we consider the case when $C=2$. Fig. 6 compares the communication cost of each algorithm to achieve different target accuracies on the MNIST and CIFAR10 datasets. Since the communication cost of HFLDD is much lower than FedAvg and FedSeq in this scenario, we use a logarithmic scale (base 10) on the y-axis. From the figure, it can be observed that as the target accuracy increases, HFLDD consistently maintains the lowest communication cost.
TABLE IV: The impact of $K$ on the test accuracy of HFLDD

### v2

@@ VI Experimental Evaluation
In this section, we first analyze the performance of the proposed HFLDD and compare it with baseline methods to demonstrate how our method can mitigate the impact of label distribution skewness. Subsequently, we compare the communication overhead of different methods. Finally, we investigate the influence of several key hyperparameters and model architectures on the performance of HFLDD.
VI-A Experimental Setup
VI-A1 Datasets
@@
Our experiments are carried out on three widely used datasets MNIST [49], CIFAR10 [50], and CINIC10 [51]. MNIST consists of 50,000 binary images of handwritten digits, which can be divided into 10 classes. CIFAR10 consists of 50,000 daily object images, which can also be divided into 10 classes. CINIC10, as an extended version of CIFAR10, comprises 270,000 images and retains the original 10 classes of CIFAR10. By incorporating downsampled images from ImageNet, CINIC10 significantly increases the complexity and diversity of the dataset. In our non-IID setting, we investigate scenarios in which each client has 1, 2, or 10 of the ten classes. In our experiments, each dataset is divided into 80% for training and 20% for testing.
@@
In addition to training and testing the local client dataset, we also need additional global datasets to generate soft labels for label knowledge collection. We randomly select 1,000 samples from other similar datasets and construct the corresponding global dataset. For the MNIST experiment, the global dataset is constructed from the FMNIST dataset [52]. For the CIFAR10 and CINIC10 experiments, the global dataset is constructed from the Tinyimagenet dataset [53].
VI-A2 Models
In our experiments, the applied models are all based on Convolutional Neural Network (CNN) [3]. For the MNIST dataset, we use a shallow CNN consisting of two convolutional layers with ReLU activations and max pooling, followed by flattening and three fully connected layers. The model contains 44,426 parameters. For the CIFAR10 dataset, a deeper CNN is adopted with four convolutional blocks. Each block includes two convolutional layers with ReLU activations, batch normalization, and a max pooling layer. After flattening, a single fully connected layer is applied. The total number of parameters is 1,020,160. For the CINIC10 dataset, the model consists of three repeated convolutional blocks, each comprising a convolutional layer with 128 filters, followed by an instance normalization layer, a ReLU activation function, and an average pooling layer. The total number of model parameters is 320,010. The input for all models is $32\times 32\times 3$ image and the output is a 10-dimensional one-hot class label. It should be noted that the proposed HFLDD is not limited by the model architectures, and it can be applied to any model. In the experiments, we also test the performance of HFLDD using different models.
@@ VI-A2 Models
(g) CINIC10 ($C=1$)
(h) CINIC10 ($C=2$)
(i) CINIC10 ($C=10$)
Figure 3: Learning curves of the global model on MNIST, CIFAR10 and CINIC10.
@@
(g) CINIC10 ($C=2$)
(h) CINIC10 ($C=2$)
(i) CINIC10 ($C=10$)
Figure 4: Heterogeneous topology formed by client clustering, where $N=100$, $K=10$. The bottom right corner of each subfigure shows the results of homogeneous clustering, where different colors represent clients from different homogeneous clusters. In each subfigure, the clients connected with the server are the clusters heads. Each cluster head and its connected clients construct a heterogeneous cluster.
VI-A3 Baselines
To evaluate the performance of our proposed HFLDD, we select five typical baseline methods for comparison:
  • FedAvg [11], a classic FL framework, where each client conducts local training and the server aggregates all the local models for global model update;
  • FedProx [11], a classic regularization-base FL framework, which adopt regularization terms to adjust local updates, and mitigate client drift on non-IID data;
  • FedSeq [34], a classic HFL framework, which utilizes client clustering and sequential training to enhance model performance;
  • DynaFed [42], a dataset distillation-based FL framework, which synthesizes distilled data by matching the training trajectory of the global model on the server, and then fine-tunes the aggregated model using the generated data
@@ VI-A4 Implementation Details
Our experiments are carried out on a machine equipped with a Tesla V100 GPU and implemented using PyTorch, we fixed the number of clients $N$ to 100. For the pre-training phase, both the MNIST, CIFAR10 and CINIC10 datasets were used with the following settings: learning rate $\eta=0.01$, local update steps in the pre-training phase $E_{1}=10$, a batch size of 64, and the SGD optimizer. These settings ensure that each client can effectively extract knowledge from its local data, thereby enhancing the discriminability for subsequent clustering tasks. To construct the heterogeneous topology, we set the clustering of K-Means with $K=10$. The size of the distilled dataset, denoted by $|\tilde{\mathcal{D}_{i}}|$, is set to 400, which means that each client generates 400 synthetic samples. Additional compression mechanisms are not applied to the distilled data beyond the dataset distillation process itself. The learning rate for optimizing the distilled data is 0.004, and the maximum number of distillation iterations is 3000. In each iteration, a minibatch of 10 samples is drawn from the original dataset to compute the kernel regression loss and update the distilled data. The regularization coefficient for kernel loss is fixed at $\lambda=1\times 10^{-6}$. Following the transmission of the distilled data, the global model training begins. Each cluster head conducts local training with $E_{2}=2$ local update rounds and a batch size of 32. The server performs global model updates for $T=300$ rounds. All other configurations are consistent with the pretraining phase.
@@
To ensure a fair comparison, FedAvg and FedSeq are configured with the same number of global update rounds and local update epochs as HFLDD. Furthermore, in the FedSeq algorithm, the number of clusters is set to 10, with each cluster containing 10 clients. For DynaFed, the server synthesizes 15 distilled data per class. The global model is then fine-tuned for 20 epochs using a batch size of 256. For centralized learning, the training batch size is 64, and the learning rate is 0.01.
VI-B Training Performance
@@
To comprehensively show the training performance of the proposed HFLDD, we compare the HFLDD with the baseline methods in both the non-IID and IID scenarios.
VI-B1 non-IID Scenario
@@
We consider two severe non-IID settings: 1) $C=1$, where each client holds only one class of data, making it the most challenging setting; 2) $C=2$, where each client holds two classes of data. The data labels among different clients are still severely unbalanced, but partially overlap.
@@
In Fig. 3, we compare the performance of the HFLDD training with the baseline methods on the datasets of MNIST, CIFAR10, and CINIC10, respectively. When $C=1$, as shown in Fig. 3a, Fig. 3d, and Fig. 3g, it can be observed that our proposed HFLDD algorithm outperforms FedAvg, FedSeq, and DynaFed in all datasets. Specifically, on the CIFAR10 dataset, HFLDD achieves an accuracy of 74.0% after 300 communication rounds, significantly outperforming FedAvg at 27.0%, FedProx at 26.3%, FedSeq at 19.6%, and DynaFed at 55.7%. In the CINIC10 dataset, HFLDD achieves an accuracy of 56.7%. Compared to FedAvg, FedProx, FedSeq, and DynaFed, it shows relative improvements of 39.3%, 57.1%, 8.8%, and 41.4%, respectively. On the MNIST dataset, although the final accuracy of HFLDD is similar to that of FedSeq, it is clear that HFLDD converges faster. DynaFed also achieves a similar final accuracy, but the use of distilled data introduces instability in the later stages of training. When $C=2$, on the CIFAR10 dataset, as shown in Figs. 3e and 3g, HFLDD achieves final test accuracies of 78.5% on the CIFAR10 dataset and 59.5% on the CINIC10 dataset, respectively, both of which outperform FedAvg, FedProx, FedSeq and DynaFed. In the MNIST data set, as shown in Fig. 3b, HFLDD, FedSeq, and DynaFed achieve nearly identical training results, and all outperform FedAvg and FedProx. Additionally, under the non-IID settings described above, the performance gap between HFLDD and CL is the smallest.
@@
Figs. 4a, 4b, 4d, 4e, 4g and 4h specifically illustrate the heterogeneous topology formed by HFLDD in the non-IID scenario. In Fig. 4d, Fig. 4g and Fig. 4h, the numbers of clients in all heterogeneous clusters are equal. In Figs. 4a, 4b, and 4e, a few isolated nodes appear in the heterogeneous topology. However, they do not significantly affect overall performance. This is because the data distribution in most heterogeneous clusters is approximately the same as the global data distribution, effectively compensating for overall performance.
@@
(c) CINIC10
Figure 5: Performance comparison of HFLDD without dataset distillation (HFLDD w/o DD) with baseline methods on three datasets when $C=10$.
VI-B2 IID Scenario
In this case, we set $C=10$, where each client holds all ten data classes, and the data labels among all clients are balanced. which significantly reduces the benefit of cross-client knowledge aggregation. On the MNIST dataset, as shown in Fig. 3c, all methods achieve nearly identical test accuracies. In the CIFAR10 and CINIC10 dataset, as shown in Fig. 3f and Fig. 3i, HFLDD exhibits slightly inferior performance compared to some baseline methods. However, it remains competitive in the IID setting.
This performance gap can be attributed to two main factors. First, HFLDD trains the global model using a combination of the local data from the cluster head and the distilled data collected from its cluster members. The dataset distillation process inevitably introduces information loss. It is important to emphasize that this information loss is not unique to the IID case but exists across all levels of data heterogeneity. In non-IID settings ($C=1$ or $C=2$), HFLDD still outperforms the baseline methods even with relatively small distilled datasets. To further investigate the impact of dataset distillation, we conduct an experiment in Section VI-D by varying the size of the distilled dataset under different levels of heterogeneity.
Second, under IID condition where each client holds data from all 10 classes, as shown in Figs. 4c, 4f and 4i, the number of clients in each heterogeneous cluster is imbalanced. After cluster heads collect the distilled data, each of them holds data that clearly include all 10 classes, meaning the label distributions across cluster heads are identical. However, due to the different number of clients in each cluster, the amount of data held by each cluster head differs, which leads to quantity skew (each cluster head contains a different number of samples). As discussed in previous work [46], such an imbalance has a minor effect on performance when weighted averaging is used during aggregation. To further assess its influence in our framework, we designed an ablation experiment in which the distilled data transmission step was replaced by direct transmission of the original local data to the cluster head. This variant, denoted by HFLDD w/o DD, removes the influence of dataset distillation. As shown in Fig. 5, HFLDD w/o DD achieves a performance comparable to or even better than the baseline methods in all three datasets, demonstrating that the impact of the cluster size imbalance caused by clustering is indeed minor.
The above analysis and experimental evidence indicate that the observed performance gap between HFLDD and other baseline methods under IID condition is mainly caused by the information loss introduced by dataset distillation, rather than the clustering mechanism.
Metrics HFLDD FedAvg FedProx FedSeq DynaFed
Accuracy=80% rounds/Accuracy 5/80.63% 17/80.54% 20/80.6% 1/84.32% 2/85.1%
Traffic (MB) 58.8M 593.15M 694.8M 39.0M 84.7M
Ratio 1X 10.1X 11.8X 0.7X 1.4X
Metrics HFLDD FedAvg FedProx FedSeq DynaFed
Accuracy=70% rounds/Accuracy 10/70.5% 35/70.3% 43/70.3% 4/71.2% 30/70.4%
Traffic (MB) 1973.7M 27630.4M 33856.9M 2296.0M 23738.7M
Ratio 1X 14.0X 17.2X 1.2X 12.0X
TABLE IV: Communication cost when achieving given accuracy on CINIC10 under IID conditions
Metrics HFLDD FedAvg FedProx FedSeq DynaFed
Accuracy=60% rounds/Accuracy 40/60.1% 44/60.0% 117/60.1% 11/60.7% 50/60.2%
Traffic (MB) 2661.4M 10864.6M 28687.4M 1745.7M 12329.5M
Ratio 1X 4.1X 10.8X 0.7X 4.6X
(a) MNIST
(b) CIFAR10
(c) CINIC10
Figure 6: Communication traffic consumption of three methods when achieving given target accuracy on MNIST, CIFAR10 and CINIC10 under non-IID conditions $(C=2)$
VI-C Communication Cost
The analytical communication cost of HFLDD is detailed in Section VI-B. In this part, we compare the communication cost of HFLDD with that of baseline methods, including FedAvg, FedProx, FedSeq and DynaFed.
The communication cost of FedAvg over $T$ rounds of training is given by Eq. (3). The communication costs of FedProx and DynaFed over $T$ rounds are the same as that of FedAvg. For FedSeq, the communication cost comes from parameter sharing between clients within cluster and across clusters, it can be expressed as
@@ VI-C Communication Cost
Considering that when the training converges, the communication rounds and test accuracy of different methods are different, it would be unfair to compare the communication cost under a given communication rounds. Therefore, we compare the number of communication rounds and total communication cost required by different methods to achieve a certain test accuracy.
In the IID scenario, the target test accuracies for the MNIST, CIFAR10, and CINIC10 datasets are set to 80%, 70%, and 60%, respectively. Table II to Table IV present the number of communication rounds and total communication cost required by different methods to reach these target accuracies. The ratio indicates the relative communication cost of each method compared to HFLDD. Experimental results show that HFLDD incurs slightly higher communication cost than FedSeq on MNIST and CINIC10, but remains lower than that of FedAvg, FedProx, and DynaFed. On CIFAR10, the communication cost of HFLDD is the lowest among all the methods.
@@
In the Non-IID scenario, when $C=1$, due to the severe imbalance of data labels, the baseline methods may not converge and it is infeasible to find proper target test accuracy that sufficiently reflects the communication cost of each method. Therefore, we consider the case when $C=2$. Fig. 6 compares the communication costs of different methods to achieve given target accuracies on the MNIST, CIFAR10 and CINIC10 datasets. Since the communication cost of HFLDD is significantly lower than that of the baseline methods in this scenario, we use a logarithmic scale (base 10) on the y-axis. From the figure, it can be observed that as the target accuracy increases, HFLDD consistently maintains the lowest communication cost.
(a) $C=1$
(b) $C=2$
(c) $C=10$
Figure 7: Learning curves of the global model with different distilled data sizes across varying data heterogeneity on CIFAR10.
VI-D Impact of distilled data size
The size of the distilled dataset significantly affects the amount of information retained from the local data of cluster members, which directly influences the performance of the global model. To investigate this impact, we evaluate the convergence behavior of HFLDD under varying levels of data heterogeneity on the CIFAR10 dataset, using different distilled data sizes. As shown in Fig. 7, increasing the distilled data size leads to notable improvements in both convergence speed and final test accuracy. This observation aligns with the theoretical expectation: a larger distilled dataset can better approximate the original data distribution, thereby reducing the distillation approximation error and improving the performance of the global model.
VI-E Impact of Network Architecture
To evaluate the generalizability and adaptability of our proposed framework across various network architectures, we conducted additional experiments using a diverse set of models on the CIFAR10 dataset, including MobileNetV2[54], ResNet18[55], and ShuffleNet[56]. These models cover a broad spectrum of architectural characteristics, such as lightweight mobile models and deeper residual networks.
Table V lists the test accuracies of different methods using the above three models under different levels of client heterogeneity. Overall, HFLDD exhibits stable and competitive performance across all model architectures and heterogeneity settings. Under severe non-IID settings, i.e., when $C=1$ and $C=2$, the test accuracies of HFLDD are largely higher than FedAvg and FedSeq, and mostly higher than that of DynaFed. Under the IID setting, i.e., when $C=10$, HFLDD remains competitive with the baseline methods, demonstrating strong adaptability across different model architectures.
TABLE V: Test accuracies of different methods using different models on CIFAR10
Method $C$ Model
MobileNetV2 ResNet18 ShuffleNet
HFLDD $C=1$ 65.8% 71.0% 58.3%
$C=2$ 71.2% 73.5% 66.1%
$C=10$ 71.5% 74.6% 71.9%
FedSeq $C=1$ 11.5% 9.8% 11.4%
$C=2$ 30.1% 27.9% 26.8%
$C=10$ 72.9% 74.9% 73.6%
FedAvg $C=1$ 11.0% 11.6% 10.3%
$C=2$ 30.2% 32.6% 45.6%
$C=10$ 68.9% 61.0% 68.5%
DynaFed $C=1$ 44.8% 55.6% 32.7%
$C=2$ 57.9% 73.7% 45.7%
$C=10$ 62.9% 78.6% 68.6%
TABLE VI: The impact of $K$ on the test accuracy of HFLDD